OpenStudy (anonymous):

The number of movie theaters in Kentucky is 22 more than twice the number of movie theaters in Ohio. If there are 154 movie theaters in Kentucky and Ohio, how many movie theaters are there in Ohio?

OpenStudy (anonymous):

Let no of movie theaters in ohio be x and movie theater in kentucky be y then y=2x+22-----(22 more than twice the number of movie theaters in Ohio) and x+y= 154---( 154 movie theaters in Kentucky and Ohio) so we can write x+2x+22=154 3x=154-22 3x=132 x=44

OpenStudy (anonymous):

ie no of theater in ohio is 44 i think you got it @Derg

OpenStudy (anonymous):

Thank you!